What is the volume of a cylinder that is 18 inches high and has a diameter of 12 inches?
Answer by nyc_function(2741)

V = pir^2h.


The diameter is given to be 12 inches. We need the radius, which is 1/2 of the diameter. The radius is 6 inches. We know the height to be 18 inches high.


Plug into the above formula and solve for V.


V = pi(6)^2(18)


V = 3.14(36)(18)


V = 2034.72 inches)^3


NOTE: If you need the answer to be in terms of pi, then do not use 3.14 for pi.


V = 648pi (inches)^3